OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (anonymous):

@e.cociuba

OpenStudy (e.cociuba):

y=-5/3x constant is -5/3 it is direct .. y = kx, k = -5/3 :))

OpenStudy (anonymous):

A. no B. yes; k = -5/3 C. yes; k = 5/3 D. yes; k = 3/5

OpenStudy (anonymous):

so b

geerky42 (geerky42):

Yeah.

OpenStudy (e.cociuba):

It's B.) :))

OpenStudy (anonymous):

Thanks Guys

OpenStudy (e.cociuba):

Np ;)